Tuesday 22nd June 2021

Infrastructure PAR: connectivity issue / high latency


We are currently having connectivity issue or high latency to some part of our Paris infrastructure. Our network provider is aware of the issue and is currently investigating.

10:03 UTC: It seems like the issue is only affecting one of the datacenter. Applications that use services deployed on another datacenter might suffer from connectivity issue or increased latency

10:15 UTC: We are removing the IPs of the affected datacenter from all DNS records of load balancers (public, internal and Clever Cloud Premium customers) and are awaiting more info from our network provider.

10:19 UTC: Packet loss and latency have been going down from 10:12 UTC and it seems to be back to normal now. We are awaiting confirmation of the actual resolution of the incident.

10:23 UTC: We are working on resolving issues caused by this network instability and making sure everything works fine.

10:25 UTC: Logs ingestion is fixed. We are working on bringing back Clever Cloud Metrics.

10:31 UTC: IPs removed from DNS records at 10:15 UTC will be added back once we have confirmation that the network issue is definitely fixed.

10:41 UTC: Full loss of connectivity between the two Paris datacenters for a few seconds around 10:39 UTC. We are still experiencing packet loss now. Our network provider is working with the affected peering network on this issue.

10:45 UTC: The two Paris datacenters are unreachable depending on your own network provider.

10:49 UTC: Network is overall very flaky. Our network provider and peering network provider are still investigating.

10:57 UTC: According to our network provider, many optical fibers in Paris are deteriorated. Some interconnection equipment might be flooded. We are waiting for more information.

11:02 UTC: (Network and infrastructure inside each datacenter are safe. The issue is clearly happening outside the datacenters.)

11:13 UTC: Network is still flaky. Overall very slow. We are still waiting for a status update from our network and peering providers.

11:20 UTC: Network seems better towards one of the datacenters. We invite you to remove all IPs starting by "46.252.181" from your DNS.

11:42 UTC: Still waiting for information from our network providers. Still no ETA.

12:16 UTC: Network loss between the datacenters has lowered a bit. Console should be more accessible.

12:21 UTC: Connections are starting to come back UP. We are still watching and waiting for more information from our network providers.

12:30 UTC: Info from provider: over the 4 optical fibers, 1 is "fine". They cannot promise this one will stay fine. They are still working on it. Teams have been dispatched on the premises.

13:15 UTC: Network is still stable. We are keeping Metrics down for now as it uses a significant amount of bandwidth between datacenters.

13:48 UTC: A second optical fiber is back UP. According to our provider, "it should be fine, now". The other two fibers are still down. The on-site teams are analysing the situation.

13:41 UTC: You can now add back these IPs to your domains:

@ 10800 IN A
@ 10800 IN A

15:35 UTC: We are bringing Clever Cloud Metrics back up. It's now ingesting accumulated data in the queue while the storage backend was down.

16:45 UTC: Clever Cloud Metrics ingestion delay is back to normal.

17:16 UTC: The situation is currently stable but may deteriorate again. We are closely monitoring it. A postmortem will be published in the following days. If the issue comes back, this incident will be updated again. Sorry for the inconvenience.

17:31 UTC: A 30 seconds network interruption happened between 17:22:42 and 17:23:10, it was an isolated maintenance event done by the datacenter's network provider.


07:01 UTC: This incident has been set to fixed as everything has been working fine, as expected, since the second optical fiber link has been restored except for the incident mentioned in the previous update. Do note that as of now we are not at the normal redundancy level as the other two optical fiber links are still down. We will update this once we have more information.

10:23 UTC: We have confirmation that a non-redundant third optical fiber link has been added at 00:30 UTC, this is only meant to add bandwidth capacity, it does not solve the redundancy issue. However, our network provider also tells us that their monitoring shows that the redundant link just came back up; although this may be temporary and the link may not be using the usual optical path.

16:13 UTC: The redundant link that came back at 10:23 UTC is stable. It may be re-routed to use another physical path at some point but we can now consider that our inter-datacenter connectivity is indeed redundant again.